OpenStudy (rock_mit182):

Rationalize the denominator and simplify. \frac{ -7 }{3sqrt{2 +1}}

OpenStudy (anonymous):

can you draw this so i can see what the problem is?

OpenStudy (rock_mit182):

\[\frac{ -7 }{3\sqrt{2 +1}}\]

OpenStudy (anonymous):

ah thank you

OpenStudy (rock_mit182):

the +1 is apart

OpenStudy (rock_mit182):

\[\frac{ 7 }{ 3\sqrt{2} +1 }\]

OpenStudy (anonymous):

oh ok

OpenStudy (rock_mit182):